OpenStudy (anonymous):

what is the slope of the line whose equation is x=-8

OpenStudy (anonymous):

y = -8x

OpenStudy (anonymous):

i haven't done this type of problem in a while...ignore me

OpenStudy (anonymous):

slope=-8

OpenStudy (anonymous):

it is vertical so it does not have a slope

OpenStudy (anonymous):

there is no slope here, and no slope does not mean the slope is zero, it means it does not exist

OpenStudy (anonymous):

what is the slope of the line whose equation is x-8

OpenStudy (anonymous):

for y=x-8 the slope is 8

OpenStudy (anonymous):

that is not an equation. it could be \[y=x-8\] or \[x=-8\] but \[x-8\] is an expression. you need an equal sign for an equation

OpenStudy (anonymous):

if it is \[y=x-8\] the slope is 1

OpenStudy (anonymous):

if it is \[x=-8\] the slope does not exist

OpenStudy (anonymous):

for y=x-8 the slope =1

OpenStudy (anonymous):

i think this equation is undefined

OpenStudy (anonymous):

please type in the whole question so we know what it is. you need an equal sign somewhere

OpenStudy (anonymous):

it's a vertical line with no intercepts

OpenStudy (anonymous):

this is the general simple formula of a line: y=mx+b, where m is the slope, b is the y- intercept

OpenStudy (anonymous):

what is the slope of the line whose equation is x=-8 this is the whole question.

OpenStudy (aroub):

the slope is undefined

OpenStudy (anonymous):

how about y=-9

OpenStudy (aroub):

this a horizontal line so i guess the slope must be zero

OpenStudy (anonymous):

how can i graph 3x-4y-1=27 using the slope of and y intercept to graph

OpenStudy (anonymous):

the slope of x=a (a=constant) would b undefined so slope of x=-8 is undefined

OpenStudy (aroub):

u want to find the x and the y intercept?

OpenStudy (anonymous):

im so thankful for the answers

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yw

OpenStudy (aroub):

ok! so if u want to find the x-intercept u have to substitute y by zero and if u want to find the y-intercept u have to substitute x by zero

OpenStudy (anonymous):

is my answer is right.. y=-3x/4+7

OpenStudy (anonymous):

hey aroub and raheen thanks for your reply

OpenStudy (aroub):

3x-4y-1=27 ---> -4y=-3x+1+27 , -4y=-3x+28 , y=3/4x-28/4

OpenStudy (anonymous):

for 3x-4y-1=27 4y=3x-28 y=(3/4) x-7

OpenStudy (aroub):

wait but u need the x and y intercepts right??

OpenStudy (anonymous):

yes this is the last question how about 6x=2y+12

OpenStudy (aroub):

wait i dont understand u wont the x-y intercepts or get it back to the form y=mx+b

OpenStudy (aroub):

??

OpenStudy (aroub):

gtg as soon as i come back! ill reply to all ur question ok? soorry bbyyeeee

OpenStudy (anonymous):

my prof said this use the slope and y intercept to graph the following equation 3x-4y-1=27 and 6x=2y+12

OpenStudy (anonymous):

6x=2y+12 y=3x-6

OpenStudy (anonymous):

hey im so thankful that u guys help through this

OpenStudy (anonymous):

yw